OpenStudy (anonymous):

If the scale factor of two similar solids is 5:13, what is the ratio their corresponding areas and volumes?

OpenStudy (anonymous):

@Michele_Laino Please help me!

OpenStudy (anonymous):

@dan815 Please help me!

OpenStudy (michele_laino):

Let's suppose that your solids are two cubes, one whose edge is L_!, and the other whose edge is L_2. Then we can write: \[\Large \frac{{L_1^3}}{{L_2^3}} = \frac{5}{{13}}\] or: \[\Large \frac{{L_1^{}}}{{L_2^{}}} = \sqrt[3]{{\frac{5}{{13}}}}\]

OpenStudy (anonymous):

I think it is 25:169 and 125:2,197. Am i correct?

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

or it is 125:2,197 and 25:169

OpenStudy (michele_laino):

no, I don't think, since the area of each cube is: \[\Large \begin{gathered} {S_1} = 6L_1^2 \hfill \\ {S_2} = 6L_2^2 \hfill \\ \end{gathered} \] then their ratio is: \[\Large \frac{{{S_1}}}{{{S_2}}} = \frac{{6L_1^2}}{{6L_2^2}} = \frac{{L_1^2}}{{L_2^2}} = {\left( {\frac{{L_1^{}}}{{L_2^{}}}} \right)^2} = {\left( {\sqrt[3]{{\frac{5}{{13}}}}} \right)^2} = {\left( {\frac{5}{{13}}} \right)^{2/3}}\]

OpenStudy (anonymous):

is it 10:26 and 15:39 i get different answers every time

OpenStudy (michele_laino):

sorry I have ,ade an error, you are right since your problem says: \[\Large \frac{{L_1^{}}}{{L_2^{}}} = \frac{5}{{13}}\] so: \[\Large \begin{gathered} \frac{{{S_1}}}{{{S_2}}} = \frac{{6L_1^2}}{{6L_2^2}} = \frac{{L_1^2}}{{L_2^2}} = {\left( {\frac{{L_1^{}}}{{L_2^{}}}} \right)^2} = {\left( {\frac{5}{{13}}} \right)^2} \hfill \\ \frac{{{V_1}}}{{{V_2}}} = \frac{{L_1^3}}{{L_2^3}} = {\left( {\frac{5}{{13}}} \right)^3} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops...I have made

OpenStudy (anonymous):

so it is 10:26 and 15:39

OpenStudy (michele_laino):

no, it is as you write above, namely: 25:169 and 125:2,197.

OpenStudy (anonymous):

Thanks. can you help with more please?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

what is the scale factor of a cube with the volume of 343ft ^3 to a cube with volume 2,744ft .^3

OpenStudy (anonymous):

i think the answer is 2:1

OpenStudy (michele_laino):

the requested scale factor, is given by the subsequent ratio: \[\Large scale\;factor = \sqrt[3]{{\frac{{343}}{{2744}}}} = ...?\]

OpenStudy (michele_laino):

so you are right: it is 1:2

OpenStudy (anonymous):

thanks. a few more?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

the volume of a sphere is 3,000 pi m ^3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (michele_laino):

the volume V of a sphere whose radius is R, is given by the subsequent formula: \[\Large V = \frac{{4\pi }}{3}{R^3}\]

OpenStudy (anonymous):

i think it's 1079 m ^2

OpenStudy (michele_laino):

so the radius is given by the inverse formula: \[\Large R = \sqrt[3]{{\frac{{3V}}{{4\pi }}}}\] and the surface is therefore: \[\Large S = 4\pi {R^2} = 4\pi {\left( {\sqrt[3]{{\frac{{3V}}{{4\pi }}}}} \right)^2}\]

OpenStudy (anonymous):

so its 2158?

OpenStudy (michele_laino):

so we have:better is 2157 meters^2

OpenStudy (anonymous):

thanks. I have another one.

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

a spherical balloon has a circumference of 21 cm . What is the approximate surface area of the balloon to the nearest square cm.

OpenStudy (anonymous):

i think it is 561 cm

OpenStudy (michele_laino):

I think that that circle is athe maximum circle, so the radius R of your balloon, is: \[\Large R = \frac{{circumference}}{{2\pi }} = \frac{{21}}{{2\pi }}\] then the requested surface, is: \[\Large S = 4\pi {R^2} = 4\pi {\left( {\frac{{21}}{{2\pi }}} \right)^2} = ...?\]

OpenStudy (anonymous):

okay so the answer is 346?

OpenStudy (michele_laino):

I got 144

OpenStudy (anonymous):

i got 140...

OpenStudy (michele_laino):

ok! that's right!

OpenStudy (anonymous):

thanks.

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

find the surface area of the sphere with the given dimension leave your answer in in terms of pi. radius of 60 m

OpenStudy (michele_laino):

The requested surface S is: \[\Large S = 4\pi {R^2} = 4\pi {\left( {60} \right)^2} = \pi \times 4 \times 3600 = ...?\]

OpenStudy (anonymous):

14, 400 pi?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thanks.

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

a rectangular pyramid fits exactly on top of a rectangular prism

OpenStudy (anonymous):

the prism has a length of 18cm width of 6 cm and height of 9 cm. the pyramid has a height of 15 cm . find the volume of the composite space figure.

OpenStudy (michele_laino):

please wait a moment, someone is calling me to the phone

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

ok! here I am

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

the volume of the prism is: \[\Large {V_{prism}} = 18 \times 6 \times 9 = ...?\] whereas the volume of pyramid is: \[\Large {V_{pyramid}} = \frac{1}{3}18 \times 6 \times 15 = ...?\]

OpenStudy (anonymous):

i think it is 1512 cm ^3?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thanks. one more then i have to go

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

the lateral area of a cone is 574 pi cm ^2 , the radius is 29 cm. what is the slant height to the nearest tenth of a cm?

OpenStudy (anonymous):

i think it is 12.6

OpenStudy (michele_laino):

the lateral surface S of a cone, is given by the subsequent formula: \[S = \frac{1}{2} \times C \times h\] where C is the circumference, and h is the slanted height, so:

OpenStudy (anonymous):

6.3?

OpenStudy (michele_laino):

\[\begin{gathered} S = \frac{1}{2} \times C \times h = \pi \times R \times h \hfill \\ \hfill \\ h = \frac{S}{{\pi R}} = \frac{{574}}{{29 \times \pi }} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

9.9? i keep getting different answers..

OpenStudy (michele_laino):

I got 6.3 cm

OpenStudy (anonymous):

that was my second answer.. but thank you so much . i will be back tonight for more help!

OpenStudy (michele_laino):

ok! Thank you!